User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting 23rd November

Chair: Arwin

Present: Patrick, Jonny, Lucia, Tiege, Liz, Ross, Mas, Bryn, Jamie, Daniel

Apologies: Jesse, Ellen


General Issues:

  • Cont17
    1. 15 day cont, 3 diff groups of scheds, all schedules will be pre dudged. Email to follow with the deets. End message and start messages need to be sent after every exp. 5 mins gap between experiments.
  • Please check your emails prior to every shift start, even if the experiment is a vanilla r1 or r4.
  • SAR tracks included in regular experiments.
    1. once every 10 days, schedules will be modified to include the satellites. Be present during these times and monitor the instruments for any errors, include missed scans in the handover notes.
  • Ke will probably have power issues as the storm season has started. Some UPSs have been fixed, hopefully we are somewhat resistant to the power surges.
  • January roster will be done before Christmas.

Issues in Handover Notes:

  • r4817: If the delay difference is massive and a reconfigure/power cycle to DBBC + fmset sync hasn't fixed it, try running counter in the eremote control to reset 'maserdelay' and 'clkoff'.
  • running counter is safe during experiments.
  • Yg drives getting stuck periodically. Happened for r1818 and r4819
  • aov018: Eremote control not connected, in such cases start the experiment using the VNC to relevant pcfs computer. To get alarms you'll need to open a terminal on ops8 and type example ./stream_log hb, repeat for any other stations needed (dont close the terminal). In the log monitor select file '/tmp/hb.log' or equivalent station, as usual run 'cloff' before reading the file in the log monitor.
  • r1819: YG error m5 -104 mk5cn: time-out, connection closed. antenna=operate and antenna=open fixed it.

Most likely a network issue or bad drives in the module? Monitor mk5 get_stats and make sure there are no other apparent issues. checkmk5 can only be called once the recording has stopped.

  • r1819: Ke drives issues, needed a power reset on site : “Turned the main switch off for 20 secs then back on. The emergency reset light was blinking and after the turn off/on the light stopped blinking.” He mentioned there were some power surges earlier on in the night, and suggested that the power may have gone off and on multiple times during the evening.
  • r4819: Ke had issues with drives and erc having the wrong time : Needed to run 'hwclock -w' again to change hardware clock, then disable fscking on reboot by editing '/etc/fstab' as root and changing the last number on the line /dev/md0 to 0 (instead of 1), then rebooting the pcfske. Now ERC and fs have the correct time.

New Scripts

  • Have a new script that setups up the mk5's of all the stations in one go.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_11_23.txt · Last modified: 2017/11/23 03:01 by Arwin Kahlon